### v1

@@ Abstract
Deep learning models have been shown to be a powerful solution for Time Series Classification (TSC). State-of-the-art architectures, while producing promising results on the UCR and the UEA archives , present a high number of trainable parameters. This can lead to long training with high CO2 emission, power consumption and possible increase in the number of FLoating-point Operation Per Second (FLOPS). In this paper, we present a new architecture for TSC, the Light Inception with boosTing tEchnique (LITE) with only $2.34\%$ of the number of parameters of the state-of-the-art InceptionTime model, while preserving performance. This architecture, with only $9,814$ trainable parameters due to the usage of DepthWise Separable Convolutions (DWSC), is boosted by three techniques: multiplexing, custom filters, and dilated convolution. The LITE architecture, trained on the UCR, is $2.78$ times faster than InceptionTime and consumes $2.79$ times less CO2 and power. To evaluate the performance of the proposed architecture on multivariate time series data, we adapt LITE to handle multivariate time series, we call this version LITEMV. To bring theory into application, we also conducted experiments using LITEMV on multivariate time series representing human rehabilitation movements, showing that LITEMV not only is the most efficient model but also the best performing for this application on the Kimore dataset, a skeleton based human rehabilitation exercises dataset. Moreover, to address the interpretability of LITEMV, we present a study using Class Activation Maps to understand the classification decision taken by the model during evaluation.
@@
This is the author’s version of an article under review in the International Journal of Data Science and Analytics.
@@ 1 Introduction
Time Series Classification (TSC) has been widely investigated by researchers in recent years. Some TSC tasks include the classification of surgical evaluation [26, 46, 16], action recognition of human motion [13, 29], cheat detection in video games [41], interpretability [52], Entomology [35], etc.  Thanks to the availability of the UCR archive [8], and the UEA archive [3], the largest archives for univariate and multivariate TSC datasets, a significant amount of work has been done in this domain. Deep learning models have been proposed in the time series context for classification [25, 27, 49, 28, 22, 39], clustering [30], averaging [47], representation learning [24, 53, 50], adversarial attacks [14, 40], etc.  Even though deep learning approaches proven to be very powerful for TSC, they present a large amount of trainable parameters, which often leads to a long training time, inference time and storage usage.
@@
Assessing the utility of LITE in the case of both univariate and multivariate time series classification datasets is essential. For this reason, we propose two versions: LITE for univariate time series and LITEMV for multivariate time series. LITEMV differs form LITE only in the first layer, where we use DWSC instead of standard convolutions. This is not done in the case of the first layer of LITE, when the input is univariate because it would lead to learn only one filter, which is not suitable for the learning task (more details on DWSC are given in Section 3.1.2.
@@ 2 Background and Related Work
Time Series Classification (TSC) was widely investigated in the literature. Some work addressed this problem using machine learning algorithms by comparing similarity metrics between the time series [2], decisions based on random forest algorithm [32], etc.  The problem of most of those algorithms is that they require huge amount of CPU time to perform their calculations, and can not be parallelized on a cluster of GPUs. For these reasons, deep learning for TSC is being used in the recent years. Even though the performance of deep learning is better than machine learning algorithms, the number of parameters to be optimized may be very high. In what follows, we first define the problem at hand, then we present some works that tackled the TSC problem using machine and deep learning techniques. Finally, we present some works that addressed the training time problem of deep learning models and the large number of parameters.
@@ 2.1 Definitions
Let x be a multivariate time series of length $L$, and let $\mathcal{D}=\{\textbf{x}_{i},y_{i}\}_{i=0}^{N-1}$ be a dataset of $N$ multivariate time series $\textbf{x}_{i}$ with their corresponding labels $y_{i}$. The goal of this work is to construct an algorithm to learn how to correctly classify each input time series to its corresponding label.
@@ 2.4 Reducing Model Complexity
Even though deep learning for TSC shown to be very effective, it still has some issues. One of these issues we address in this work is the large number of parameters to be optimized, which increases the training time as well. Recently, in [9] a new approach was proposed, called ROCKET, that also includes convolution operations, but is way faster than InceptionTime. They proposed not to learn few filters with a backpropagation algorithm, but instead randomly generate a large number of filters with different characteristics. These characteristics include the length of the filter, the bias value, the dilation rate, etc.  It has been shown that on the UCR archive, no statistical significant difference can be found between InceptionTime and ROCKET. The main advantage of ROCKET compared to InceptionTime is the training and inference time. Some adaptations of ROCKET were proposed in order to boost its performance even more such as MiniROCKET [10] and MultiROCKET [45]. Knowledge distillation [20] was also approached for the TSC model called FCN [1]. In this study, the authors proposed a smaller variant of FCN with a lower number of convolution layers and filters to learn.
@@ 3.1.2 DepthWise Separable Convolution DWSC
In standard convolutions, if the input sample of length $L$ has $C_{in}$ channels and the desired output is a space with $C_{out}$ channels using a kernel of size $k$, then the number of learned parameters is $C_{in}*C_{out}*k$. The number of multiplications is $L*C{out}*C_{in}*k$.
@@ 3.2 Boosting Techniques
Dilated convolutions were not very explored for deep supervised learning on TSC but they were used in self-supervised models showing to be very effective [50]. Dilation in convolutions filters defines the number of empty cells in the kernel. Suppose a kernel of length $3$ has the following parameters $k=[k_{0},k_{1},k_{2}]$, a dilation of rate 2 will produce the following kernel $k=[k_{0},skip,k_{1},skip,k_{2}]$. The $skip$ parameter indicates that the convolution layer will not use the values of the input aligned with this index of the kernel. A visualization of the dilation effect on convolution can be seen in Figure 4. Dilation will help increasing the receptive field of a model without having to add deeper layers because the dilated kernel will find the deeper combinations in the same layer.
@@ 4.2 Implementation Details
4.3 Results and Discussion
First, the table shows that the smallest model in terms of number of parameters is the LITE with $9,814$ parameters. This is mainly due to the usage of DWSC instead of standard ones. Compared to FCN ResNet and Inception, LITE has only $3.7\%$ $1.95\%$ and $2.34\%$ of their number of parameters, respectively. Second, the fastest model in the training phase is LITE, with a training time of $0.62$ days. LITE is $2.79$, $3.08$ and $2.71$ time faster than FCN, ResNet and Inception, respectively. Third, LITE is the model that consumes the smallest amount of CO2 and Power, $0.1048$ grams and $0.2468$ Watts, respectively. Compared to the other approaches, LITE presents the fastest and most ecologic model for TSC compared to FCN, ResNet and Inception. We believe, given the factors explained above, that LITE can be used for the deployment of deep learning for TSC in small machine such as mobile phones.
4.3.2 Accuracy Performance on the UCR Archive
On the one hand, the results presented in Figure 6 show that LITE beats FCN significantly (low $P$-Value), and is statistically not significant than ResNet (high $P$-Value). The results compared to ResNet are impressive given the small complexity of LITE (almost $1.95\%$ of ResNet’s number of parameters). On the other hand, the comparison shows that LITE still is not significantly close to Inception. To study more the reason why LITE performs not as good as Inception with a large margin (more than $10\%$), we presented some characteristics of those datasets in Table 2. This table shows that some of the datasets have either long time series or small training set. Firstly, this indicates that Inception is better than LITE on long time series given its large receptive field (deeper architecture). Secondly, this points out that Inception generalizes better better in the case where the dataset presents a small training set so it can generalize better.
Furthermore, a new Multi-Comparison Matrix (MCM) evaluation tool was proposed that is stable to the variation of the addition and removal of classifiers [23]. The MCM is presented in Figures 8 and 9 to compare LITE and LITETime, respectively, to other approaches in the literature. The MCM uses the average accuracy on the UCR as an ordering metric instead of the average rank. As presented in the MCMs, LITE and LITETime perform better than FCN and ResNet on the average accuracy and is closer to the performance of InceptionTime, which is not significantly different than LITETime (high p-value). MCM has an advantage over the usage of the CD Diagram of being stable with the addition and removal of classifiers, given that the average performance would not change in this scenario unlike the average rank. Another advantage is not using a multiple test correction for the $P$-Value significance test.

### v2

@@ Abstract
Deep learning models have been shown to be a powerful solution for Time Series Classification (TSC). State-of-the-art architectures, while producing promising results on the UCR and the UEA archives , present a high number of trainable parameters. This can lead to long training with high CO2 emission, power consumption and possible increase in the number of FLoating-point Operation Per Second (FLOPS). In this paper, we present a new architecture for TSC, the Light Inception with boosTing tEchnique (LITE) with only $2.34\%$ of the number of parameters of the state-of-the-art InceptionTime model, while preserving performance. This architecture, with only $9,814$ trainable parameters due to the usage of DepthWise Separable Convolutions (DWSC), is boosted by three techniques: multiplexing, custom filters, and dilated convolution. The LITE architecture, trained on the UCR, is $2.78$ times faster than InceptionTime and consumes $2.79$ times less CO2 and power, while achieving an average accuracy of $84.62\%$ compared to $84.91\%$ with InceptionTime. To evaluate the performance of the proposed architecture on multivariate time series data, we adapt LITE to handle multivariate time series, we call this version LITEMV. To bring theory into application, we also conducted experiments using LITEMV on multivariate time series representing human rehabilitation movements, showing that LITEMV not only is the most efficient model but also the best performing for this application on the Kimore dataset, a skeleton based human rehabilitation exercises dataset. Moreover, to address the interpretability of LITEMV, we present a study using Class Activation Maps to understand the classification decision taken by the model during evaluation.
@@
This is the author’s version of an article published in the International Journal of Data Science and Analytics. The final authenticated version is available online at: https://doi.org/10.1007/s41060-024-00708-5
@@ 1 Introduction
Time Series Classification (TSC) has been widely investigated by researchers in recent years. Some TSC tasks include the classification of surgical evaluation [29, 52, 18], action recognition of human motion [14, 32], cheat detection in video games [44], interpretability [58], Entomology [38], etc.  Thanks to the availability of the UCR archive [9], and the UEA archive [3], the largest archives for univariate and multivariate TSC datasets, a significant amount of work has been done in this domain. Deep learning models have been proposed in the time series context for classification [30, 55, 31, 24, 42], clustering [33], averaging [53], representation learning [27, 60, 56], adversarial attacks [16, 43], etc.  Even though deep learning approaches proven to be very powerful for TSC, they present a large amount of trainable parameters, which often leads to a long training time, inference time and storage usage.
@@
Assessing the utility of LITE in the case of both univariate and multivariate time series classification datasets is essential. For this reason, we propose two versions: LITE for univariate time series and LITEMV for multivariate time series. LITEMV differs form LITE only in the first layer, where we use DWSC instead of standard convolutions. This is not done in the case of the first layer of LITE, when the input is univariate because it would lead to learn only one filter, which is not suitable for the learning task (more details on DWSC are given in Section 3.1.2).
@@
We note that this manuscript extends our previous work titled“”LITE: Light Inception with boosTing Techniques for Time Series Classification” [26]. We have expanded our research to include the following enhancements:
  • A comprehensive analysis on the significant impact of the number of LITE models in the LITETime ensemble;
  • An extension of the LITE architecture to handle multivariate time series data;
  • A detailed study on a real world application of human motion rehabilitation, complemented by explainability experiments;
@@ 2 Background and Related Work
Time Series Classification (TSC) is widely investigated in the literature. Some work addressed this problem using machine learning algorithms by comparing similarity metrics between the time series [2], decisions based on random forest algorithm [35], etc.  The problem of most of those algorithms is that they require huge amount of CPU time to perform their calculations, and can not be parallelized on a cluster of GPUs. However, we acknowledge that libraries such as cuML [45] do exist, which allow for efficient execution of some machine learning workloads on GPUs. For these reasons, deep learning for TSC is being used in the recent years. Even though the accuracy and efficiency of deep learning often exceed those of traditional machine learning algorithms in many contexts, the number of parameters that need to be optimized can be significantly higher. In what follows, we first define the problem at hand, then we present some works that tackled the TSC problem using machine and deep learning techniques. Finally, we present some works that addressed the training time problem of deep learning models and the large number of parameters.
@@ 2.1 Definitions
  • Let $\textbf{x}=\{x_{1},x_{2},\ldots,x_{L}\}$ denote a univariate time series of length $L$, consisting of a sequence of data points that are equally spaced in time and exhibit temporal dependencies;
  • Let $\textbf{X}=\{\textbf{x}_{1},\textbf{x}_{2},\ldots,\textbf{x}_{M}\}$ be a multivariate time series of $M$ variables, where each variable is itself a univariate time series of length $L$
  • Let $\mathcal{D}=\{\textbf{X}_{i},y_{i}\}_{i=0}^{N-1}$ be a dataset of $N$ multivariate time series $\textbf{X}_{i}$ with their corresponding labels $y_{i}$.
The goal of this work is to construct an algorithm to learn how to correctly classify each input time series to its corresponding label.
@@ 2.3 Deep Learning for TSC
Recurrent Neural Networks (RNNs), including their variants LSTM and GRU, have been investigated for Time Series Classification (TSC) [51, 30]. While these models excel in sequential data processing, they often fail to perform optimally in TSC scenarios due to their design, which does not favor the discriminative identification of complex patterns across time series. This inherent limitation makes them less suited for tasks where precise pattern recognition in time series data is crucial.
Transfer learning and few-shot learning techniques [15, 28] can be employed to reduce the training time for models like Inception by leveraging pre-trained networks. While these approaches decrease the duration of training significantly, they do not alter the model’s complexity or reduce the number of parameters, which remains a consideration for inference tasks.
@@ 2.4 Reducing Model Complexity
While deep learning for Time Series Classification (TSC) has proven effective, it faces challenges, such as the extensive number of parameters requiring optimization which can prolong training time. An alternative approach introduced by [10] is ROCKET, which utilizes convolution operations but differs significantly from typical deep learning methods like InceptionTime [31]. Unlike traditional methods that learn filters through backpropagation, ROCKET employs a strategy of generating a large set ($10,000$) of 1D convolution filters randomly following a Gaussian distribution. These filters vary in length, bias value, padding, and dilation rate, and are not learned but are fixed upon initialization. Each convolution output is then processed to extract two features: the Proportion of Positive Values (PPV) and the maximum value. This results in a substantial feature vector, creating $20,000$ features per series, which are used to train a Ridge classifier. Studies on the UCR archive [9] indicate no statistical significant difference in accuracy between InceptionTime and ROCKET, yet ROCKET offers a substantial advantage in terms of both training and inference speed.
Some adaptations of ROCKET were proposed in order to boost its performance even more such as MiniROCKET [11] and MultiROCKET [50]. Knowledge distillation [22] was also approached for the TSC model called FCN [1]. In this study, the authors proposed a smaller variant of FCN with a lower number of convolution layers and filters to learn.
@@ 3.1.2 DepthWise Separable Convolution DWSC
In standard convolutions, if the input sample of length $L$ has $C_{in}$ channels and the desired output is a space with $C_{out}$ channels using a kernel of size $k$, then the number of learned parameters is $C_{in}*C_{out}*k$. The number of multiplications is $L*C{out}*C_{in}*k$. This assumes, as with other parts of this section, that the convolution layer uses “same” padding, thereby preserving the length from the input series to the output.
@@
Equation 3 below presents the overall formulation of a DWSC layer applied on an input multivariate time series $\textbf{X}=\{\textbf{x}_{1},\textbf{x}_{2},\ldots,\textbf{x}_{M}\}$ with $M$ variables of length $L$, with $C_{out}$ as a target output dimension using a kernel size of length $k$.
$\textbf{O}=concat(\{\sum_{m=1}^{M}concat(\{\textbf{x}_{m}*\textbf{w}_{m}\}_{m=1}^{M})^{m}.w_{m,j}\}_{j=1}^{C_{out}})$ (3)
where $\textbf{w}_{m}$ is the $m^{th}$ convolution filter applied in phase 1 (DepthWise convolution), $concat(.)$ is the concatenation operation over the spatial dimension, * is the one dimensional convolution operation and $w_{m,j}$ is one trainable scalar of phase 2 (PointWise convolution) where $m~{}\in~{}[1,M]$ and $i~{}\in~{}[1,C_{out}]$, making the output of the convolution O a multivariate time series of $C_{out}$ dimensions of length $L$ assuming a padding “same”.
@@ 3.2 Boosting Techniques
Dilated convolutions were not very explored for deep supervised learning on TSC when evaluating on the UCR archive [9] but they were used in self-supervised models showing to be very effective [56]. Additionally, outside the score of the UCR archive [9], some research work have addressed the task of deep supervised learning on TSC using dilated convolutions [46, 59, 4]. Dilation in convolutions filters defines the number of empty cells in the kernel. Suppose a kernel of length $3$ has the following parameters $k=[k_{0},k_{1},k_{2}]$, a dilation of rate 2 will produce the following kernel $k=[k_{0},skip,k_{1},skip,k_{2}]$. The $skip$ parameter indicates that the convolution layer will not use the values of the input aligned with this index of the kernel. A visualization of the dilation effect on convolution can be seen in Figure 4. Dilation will help increasing the receptive field of a model without having to add deeper layers because the dilated kernel will find the deeper combinations in the same layer.
@@ 4.2 Implementation Details
4.3 Results and Discussion
First, the table shows that the smallest model in terms of number of parameters is the LITE with $9,814$ parameters. This is mainly due to the usage of DWSC instead of standard ones. Compared to FCN ResNet and Inception, LITE has only $3.7\%$ $1.95\%$ and $2.34\%$ of their number of parameters, respectively. Second, the fastest model in the training phase is LITE, with a training time of $0.62$ days. LITE is $2.79$, $3.08$ and $2.71$ time faster than FCN, ResNet and Inception, respectively. Third, LITE is the model that consumes the smallest amount of CO2 and Power, $0.1048$ grams and $0.2468$ Watts, respectively. Compared to the other approaches, LITE presents the fastest and most ecologic model for TSC compared to FCN, ResNet and Inception. We believe, given the factors explained above, that LITE can be used for the deployment of deep learning for TSC in small machine such as mobile phones. As shown in Table 1, the LITE model not only has the smallest number of parameters among the evaluated models but also has the second fastest inference time, making it well-suited for environments where model size and speed are crucial.
Although DTW is a relevant technique in time series analysis, it was not included in our comparative efficiency analysis due to its inability to be parallelized on GPUs, a feature that our proposed deep learning approaches exploit to achieve scalability and speed. Additionally, ROCKET was excluded from our efficiency comparative analysis because it is primarily implemented for multi-threaded CPU use, not GPU acceleration, which differs from the hardware focus of our study.
4.3.2 Accuracy Performance on the UCR Archive
@@
On the one hand, the results presented in Figure 6 show that LITE beats FCN significantly (low $P$-Value), and is statistically not significant than ResNet (high $P$-Value). The results compared to ResNet are impressive given the small complexity of LITE (almost $1.95\%$ of ResNet’s number of parameters). On the other hand, the comparison shows that LITE still is not significantly close to Inception. To study more the reason why LITE performs not as good as Inception with a large margin (more than $10\%$), we presented some characteristics of those datasets in Table 2. This table shows that some of the datasets have either long time series or small training set. Firstly, this indicates that Inception is better than LITE on long time series given its large receptive field (deeper architecture). Secondly, this demonstrates that the Inception model generally exhibits stronger generalization capabilities, particularly when the dataset has a smaller training set.
Figure 11: A collection of four Pareto plots showcasing the average accuracy of the four models ResNet [55] Inception [31], FCN [55] and ours LITE, over the UCR archive datasets [9] in function of the number of parameters, number of flops, training time (in days) and inference time (in days).
Furthermore, a new Multi-Comparison Matrix (MCM) evaluation tool was proposed that is stable to the variation of the addition and removal of classifiers [25]. The MCM is presented in Figures 9 and 10 to compare LITE and LITETime, respectively, to other approaches in the literature. The MCM uses the average accuracy on the UCR as an ordering metric instead of the average rank. As presented in the MCMs, LITE and LITETime perform better than FCN and ResNet on the average accuracy and is closer to the performance of InceptionTime, which is not significantly different than LITETime (high p-value). MCM has an advantage over the usage of the CD Diagram of being stable with the addition and removal of classifiers, given that the average performance would not change in this scenario unlike the average rank. Another advantage is not using a multiple test correction for the $P$-Value significance test.
As depicted in the MCM in Figure 10, LITETime outperforms Inception with statistical significance in terms of accuracy. Regarding latency, although a single LITE model operates at approximately half the inference time of Inception, the combined inference time of the LITETime ensemble, which includes five LITE models, is about 2.5 times that of Inception. However, it is crucial to note that the ensemble’s overall complexity in terms of the number of parameters is significantly reduced—eight times smaller than Inception. This substantial reduction in model complexity, coupled with the potential for parallelizing these models, suggests that improvements in parallel processing could effectively diminish the apparent latency disadvantage, offering a practical solution for scenarios where model efficiency and reduced hardware demand are critical.
In addition to Table 1, Figure 11 presents similar metrics, including a collection of Pareto plots comparing the average performance of LITE and three comoetitotrs in function of number of parameters, number of FLOPS, training time and infernece time. The plot illustrates that LITE performs as good as state-of-the-art models, such as Inception with a difference in average accuracy of only $0.89\%$, while being $42$ times smaller than Inception, having $40$ times less FLOPS, and being $2.7$ and $1.8$ times faster in terms of training and inference time respectively.